OpenStudy (anonymous):

Find the value of the sec 19° using your calculator. Answer 0.989 0.946 1.058 1.011

OpenStudy (amistre64):

my calculator only does division by primes ....

OpenStudy (amistre64):

\[cos(x)=1-\frac{1}{2!}x^2+\frac{1}{4!}x^4-\frac{1}{6!}+x^6-\frac{1}{8!}x^8...\] but youd have to convertdegrees into radians; 19*pi/180 and us it for x then flip the results to get a sec

OpenStudy (amistre64):

opps, got a typo :) \[cos(x)=1-\frac{1}{2!}x^2+\frac{1}{4!}x^4-\frac{1}{6!}x^6+\frac{1}{8!}x^8...\] thats better